High Desert Potteryworks makes a variety of pottery products that it sells to retailers. The company uses a job-order costing system in which departmental predetermined overhead rates are used to apply manufacturing overhead cost to jobs. The predetermined overhead rate in the Molding Department is based on machine-hours, and the rate in the Painting Department is based on direct labor-hours. At the beginning of the year, the company provided the following estimates:

Department
Molding Painting
Direct labor-hours 35,000 57,800
Machine-hours 89,000 36,000
Fixed manufacturing overhead cost $ 249,200 $ 560,660
Variable manufacturing overhead per machine-hour $ 2.40 -
Variable manufacturing overhead per direct labor-hour - $ 4.40

Job 205 was started on August 1 and completed on August 10. The company's cost records show the following information concerning the job:

Department
Molding Painting
Direct labor-hours 78 133
Machine-hours 370 65
Direct materials $ 942 $ 1,140
Direct labor cost $ 730 $ 1,010

Required:

1. Compute the predetermined overhead rates used in the Molding Department and the Painting Department.

2. Compute the total overhead cost applied to Job 205.

3-a. What would be the total manufacturing cost recorded for Job 205?

3-b. If the job contained 28 units, what would be the unit product cost?
Answer #1

Answers

  • [1]

Working

Molding

Painting

A [given]

Total Fixed manufacturing Overhead cost

$249,200

$560,660

Allocation base

machine hours

direct labor hours

B [given]

Amount of allocation base

                        89,000

                                     57,800

C = A/B

Fixed Overhead rate

$                         2.80

$                                      9.70

D [given]

Variable overhead rate

$                         2.40

$                                      4.40

E = C+D

Predetermined Overhead rates = ANSWER

$                         5.20

$                                    14.10

  • [2]

Total Overhead cost applied = (370 MHs x $ 5.20) + (133 DLHs x $ 14.10)
= 1924 + 1875
= $ 3799

  • [3]

Molding

Painting

Total

Direct Material Cost

$942

$1,140

$2,082

Direct Labor Cost

$730

$1,010

$1,740

Manufacturing Overhead cost applied

$1,924

$1,875

$3,799

Total manufacturing cost assigned

$3,596

$4,025

$7,621

  • [4]

Unit Product cost = $ 7621 / 28 units
= $ 272.18 per unit